OpenStudy (anonymous):

dx/dt=9-4x^2, x(0)=0 Seperate variables use partial fractions to solve the initial value

OpenStudy (anonymous):

hmm..maybe\[\frac{dx}{dt} =9-4x^2 => \frac{dt}{dx} =\frac{1}{9-4x^{2}} dx => t=\int\limits\frac{1}{9-4x^{2}} dx\]not sure tho..havent done these in a while.

OpenStudy (anonymous):

yes set up is \[\int\limits_{?}^{?}9-4x^2= \int\limits_{?}^{?}dt\] 9-4x^2 dx=\int\limits_{?}^{?} dt\]

OpenStudy (anonymous):

\[x=\int\limits9-4x^{2} dx\]?? doesnt look right tho

OpenStudy (anonymous):

right side is t + C but left side i'm lost

OpenStudy (anonymous):

sorry left side woud be interval 1/(9-4x^2) dx

OpenStudy (anonymous):

so what i had; is right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok..use trig sub

OpenStudy (anonymous):

help

OpenStudy (anonymous):

\[x=\frac{9}{4} \sin\theta \] for trig sub or you can use table of standard integrals

OpenStudy (anonymous):

need help with ∫9−4x 2 dx do i use u=9-4X^2 so du= -8x but them what?

OpenStudy (accessdenied):

I would separate the variables like so: \[ \frac{dx}{dt} = 9 - 4x^2 \\ \frac{1}{9 - 4x^2} \frac{dx}{dt} = 1 \\ \frac{1}{9 - 4x^2} dx = dt \\ \int \frac{1}{9 - 4x^2} dx = \int 1 dt \\ \int \frac{1}{9 - 4x^2} dx = t + C \] \[ \int \frac{1}{9 - 4x^2} dx \] Then, use partial fractions on this integral here.

OpenStudy (anonymous):

you can use trig-sub; partial fractions is too tedious.

OpenStudy (accessdenied):

Well, that may be true, but the original post seems to ask specifically for partial fractions... ``` Seperate variables use partial fractions to solve the initial value ```

OpenStudy (anonymous):

i need help with the partial fraction of 1/(9-4x^2) can you show me the steps, please, :-)

OpenStudy (accessdenied):

Well, the first step here is to identify the factors of the denominator. Do you know how to factor this denominator?

OpenStudy (anonymous):

yes 3-2x and 3+2x

OpenStudy (accessdenied):

Okay, good. :) We can rewrite this integral a bit then: \[ \int \frac{1}{(3 - 2x)(3 + 2x)} dx \] Partial fractions is basically the process of writing this fraction in terms of a sum of, usually, more simple fractions, like this: There are some differences depending on how the denominator is factored, but here is the specific way for our denominator: non-repeated linear factors only: \[ \frac{1}{(3 - 2x)(3 + 2x)} = \frac{A}{3 - 2x} + \frac{B}{3 + 2x} \] We just have to find the A and B here.

OpenStudy (anonymous):

ok let me try

OpenStudy (accessdenied):

I recall there are multiple ways to do this, so I will be using one specific technique that I am familiar with. (It may not be the same as what you use) Multiplying both sides by the denominator of the left-hand side, it becomes: \[ \frac{\cancel{(3 - 2x)(3 + 2x)}}{\cancel{(3 - 2x)(3 + 2x)}} = \frac{A\cancel{(3 - 2x})(3 + 2x)}{\cancel{3 - 2x}} + \frac{B(3 - 2x)\cancel{(3 + 2x)}}{\cancel{3 + 2x}} \\ 1 = A(3 + 2x) + B(3 - 2x)\\ 1 = 3A + 2Ax + 3B - 2Bx \\ 1 = (2A - 2B)x + (3A + 3B)\] Basically, we can equate the coefficients here. \(2A - 2B = 0\) \(3A + 3B = 1\) This is a system of linear equations, which may be solved for the A and B.

OpenStudy (accessdenied):

* If it is unclear where the \(2A - 2B = 0\) part comes from, remember that \(1 = 0x + 1\), and we equate the (2A + 2B) coefficient on the right side to this 0.

OpenStudy (anonymous):

the book says 1/6(3-2x)+1/6(3+2x) i dont get it

OpenStudy (anonymous):

thats the partial fraction of 1 (3−2x)(3+2x)

OpenStudy (accessdenied):

Have you read my above explanation or do not understand it?

OpenStudy (anonymous):

do not understand it because its been 22 years since I was in college and Im a litttle lost right now.

OpenStudy (accessdenied):

Sorry, was trying to make this: http://i49.tinypic.com/2hfigr8.gif Maybe this is easier to follow? It goes through the basic process to get the system, putting up the next step every 2 seconds. It stops at the end.

OpenStudy (anonymous):

i follow this p://i49.tinypic.com/2hfigr8.gif find but that still doesnt get me to 1/(6(3-2x))+1/(6(3+2x))

OpenStudy (accessdenied):

Okay, now we'll solve that linear system, which is just Algebra work now: 2A - 2B = 0 Multiply both sides by 3 3A + 3B = 1 Multiply both sides by 2 6A - 6B = 0 6A + 6B = 2 Add together to eliminate B: 12A = 2 A = 1/6 6(1/6) - 6B = 0 1 - 6B = 0 1 = 6B 1/6 = B We can put these in for A and B to get what you posted. \[ \frac{1}{(3 - 2x)(3 + 2x)} = \frac{1 \over 6}{3 - 2x} + \frac{1 \over 6}{3 + 2x} \\ \frac{1}{(3 - 2x)(3 + 2x)} = \frac{1}{6}\frac{1}{3 - 2x}\ + \frac{1}{6} \frac{1}{3x + 2}] \]

OpenStudy (accessdenied):

Now, this allows us to rewrite that integrand into a sum of two fractions. I'll factor out the 1/6 and we get: \[ \int \frac{1}{9 - 4x^2} dx = \int \frac{1}{6} \left( \frac{1}{3 - 2x} + \frac{1}{3 + 2x} \right) dx \]

OpenStudy (accessdenied):

Sorry, it's kind of late, I'll try to hurry up: We may bring the 1/6 out front and integrate each term: \[ \frac{1}{6} ( \int \frac{1}{3-2x} dx + \int \frac{1}{3 + 2x} dx ) \] From this, we can make some u-substitutions and get a form much more like: \[ \int \frac{1}{u} du \] Which is just \(\ln |u|\)

OpenStudy (accessdenied):

\[ \int \frac{1}{3 - 2x} dx\] Let \(u = 3 - 2x\) \(du = -2 dx \implies \neg \frac{1}{2} du = dx \) \[ \int \frac{1}{u} (\neg \frac{1}{2} du) \\ \neg \frac{1}{2} \int \frac{1}{u} du \\ \neg \frac{1}{2} \ln |u| \\ \neg \frac{1}{2} \ln |3 - 2x| \] I assume you are familiar with this process?

OpenStudy (anonymous):

ok i get this now... so the I get -1/12 ln (3-2x)+1/12 ln(3+2x)=t solve for x need help with this part. and yes I familiar with u sub.

OpenStudy (accessdenied):

Are you required to solve for x? Seems pretty difficult to solve this without maybe a hyperbolic trig function, and would be easier to leave in implicit form...

OpenStudy (anonymous):

maybe not i'll fid out.. thanks for your help. you were very helpful

OpenStudy (accessdenied):

The immediate process to solve is just to take x = 0, t = 0 and solve for the constant C: -1/12 ln (3-2x)+1/12 ln(3+2x)=t + C > The + C is from indefinite integration.

OpenStudy (accessdenied):

Oh, I got it. You need a bit of manipulation to do it: We could use a logarithm property to get one logarithm log(a) + log(b) = log(ab) log(a) - log(b) = log(a/b) -1/12 ln (3-2x)+1/12 ln(3+2x)=t + C (don't forget that +C since its indefinite integral) factor out a 1/12 1/12(-ln(3 - 2x) + ln(3 + 2x)) = t + C reordering 1/12(ln(3 + 2x) - ln(3 - 2x)) = t + C use log property 1/12 ln( (3+ 2x)/(3 - 2x) ) = t + C multiply both sides by 12 ln( (3 + 2x)/(3 - 2x) ) = 12t + C raise both sides as a power of e (3 + 2x)/(3 - 2x) = e^(12t+C) Multiply (3 - 2x) denominator to both sides 3 + 2x = (3 - 2x) e^(12t + C) Distribute 3 + 2x = 3 e^(12t + C) - 2x e^(12t + C) Get terms with x on one side 2x + 2xe^(12t + C) = e^(12t + C) - 3 Factor out the x. x(2 + 2e^(12t + C) = e^(12t + C) - 3 Divide by (2 + 2e^(12t + C)) x = (e^(12t + C) - 3) / (2e^(12t + C) + 2)

OpenStudy (accessdenied):

In that form, C is a bit tougher to solve for though for the initial value problem. :P

OpenStudy (anonymous):

\[x=\frac{ -2 }{ 4 }k e^{t}\] is this the correct answer for \[\frac{ dx }{ dt}=4x-2\]

OpenStudy (anonymous):

\[\frac{ dy }{ dt}=-3y-2 ; y(2) = 0 .. \] when i am solving this equation i have got \[y=\frac{ -2 }{ 3 }+Ke ^{-3t}\] .. now please help me for next step with y(2) = 0...